# Following the Human Thread in Social Navigation

### v1

@@ 4 Results
Section 4.1 presents our findings regarding Social Navigation, whereas Section 4.2 conducts an ablation analysis concentrating on adaptable information and the role of the trajectory history length. Furthermore, Section 4.3 provides a qualitative exploration of our outcomes.
@@
We utilize our SDA on Habitat 3.0[36], a simulation platform designed for human-robot interaction within domestic settings. This platform offers precise humanoid simulation capabilities with a strong focus on collaborative tasks such as Social Navigation and Social Rearrangement. It offers a vast library of avatars featuring multiple gender representations, body shapes, and appearances, facilitated by SMPL-X [34]. Furthermore, it employs an oracle policy to generate movement and behavior, enabling programmable control of avatars for navigation, object interaction, and a range of other movements.
@@
The baselines we employ in our study are drawn from Habitat 3.0 [36] and consist of:
  • Heuristic Expert: a heuristic baseline equipped with access to the environment map, employing a shortest path planner to devise a route to the current location of the humanoid. The heuristic expert operates on the following principles: When the agent is beyond a distance of 1.5 meters from the humanoid, it employs a “find” behavior, utilizing a path planner to approach the humanoid. Conversely, if the humanoid is within 1.5 meters, the expert executes a backup motion to prevent collision with the humanoid.
  • Baseline: the current SoA method [36], a recurrent neural network policy trained with DD-PPO [48], operates on a “sensors-to-action” paradigm. Inputs to this policy consist of an egocentric arm and stereo depth sensors, a humanoid detector, and humanoid GPS coordinates, while the outputs are velocity commands in the robot’s local frame. Notably, this policy operates without access to an environment map. Table 1 compares our model against the most realistic configuration, wherein the humanoid GPS data is unavailable.
  • Proximity tasks: we also adapted the Proximity Tasks defined in [9] and applied them to the baseline [36]. These tasks were proposed for a different setup of SocialNav, where the agent acts in an environment with multiple humanoids and must navigate from point A to point B while avoiding collisions. We adapted the risk and compass proximity tasks to the SocialNav setting addressed in this article. In this context, the risk has a low value (close to 0) when the agent is within 1 to 2 meters from the humanoid and a value close to 1 when the distance is less than 1 meter or greater than 2 meters. Similarly, given the presence of only one humanoid in the environment, the compass was redefined to predict the angle between the humanoid and the agent. This adjustment aims to assist during the following phase, enabling the agent to follow the humanoid while maintaining a safe distance and staying aligned with the human. The proximity tasks are jointly trained with the policy and detached during evaluation.
@@
We used the metrics for the SocialNav task as defined in [36], briefly summarized in the following. 1. Finding Success ($S$)is the ratio of the episodes where the agent successfully located and reached the human. 2. Finding Success Weighted by Path Steps ($SPS$)measures the optimality of the path taken by the agent w.r.t. the optimal number of steps needed to reach the human. 3. Following Rate ($F$)is the ratio of steps during which the robot maintains a distance of 1-2 meters from the humanoid while facing towards it relatively to the maximum possible following steps. 4. Collision Rate ($CR$)is the ratio of the episodes that ended with the robot colliding with the humanoid. 5. Episode Success ($ES$)measures the ratio of the episodes where the agent found the human and followed it for the required number of steps, maintaining a safe distance in the 1-2 meters range.
@@
In our research, the privileged information under consideration includes humanoid GPS (hGPS) and human trajectories (traj.). Humanoid GPS is represented in polar coordinates, a method of specifying a point’s position in a plane using two parameters: the distance from the point to the origin (radius) and the angle formed between a reference direction (typically the positive x-axis) and a line connecting the origin to the point (polar angle or azimuth). In our context, the origin is defined as the robot’s position; thus, its position is implicitly known along with that of the human. Conversely, trajectories solely consist of information derived from the human’s position within the environment.
@@ 4.1 Quantitative Results
In Table 1, we present the results obtained in Habitat 3.0 [36] for the Social Navigation task. The table is divided into three sections. The first one displays outcomes achieved by utilizing a heuristic expert endowed with extensive information, including its position, map data, and the humanoid’s position, granting it a competitive advantage over other methods. The subsequent section features models trained and tested using ground truth (GT) data (Baselines and Stage 1), thus establishing an upper limit for techniques utilizing more practical inputs feasible in real-world scenarios. Lastly, the final rows delineate results from methods conducting inference without privileged information. The models that use privileged information (GT) show that S1 has lower performance than the two Baselines, especially in $S$ and $SPS$. This is to be expected, considering that while trajectories solely depict the movement of the human in the environment, hGPS furnishes crucial details on locating the humanoid, offering insights into the distance and angle between them.
In the second stage, despite the absence of human trajectory input, SDA maintains the performance level of Stage 1 by adapting to the social dynamics. Our model outperforms the best techniques in the find task, increasing $S$ and $SPS$ by $6\%$.
Our approach generally improves performance in the episode success ($ES$) metric, which occurs when the agent finds the humanoid and follows it for $400$ steps. However, we emphasize that the test-time episode ends at $1500$ steps or when the agent collides with the humanoid [36], not after the $400$ follow steps. In this context, $S$, $SPS$, and $F$ metrics demonstrate how SDA, compared to the Baseline, more frequently locates the humanoid, follows a more optimal path (on average $438$ vs. $540$ steps), and follows it for a longer duration ($390$ vs. $218$ steps). Therefore, as the agent follows the humanoid longer, it has more chances to collide with it, given that the episode does not necessarily end after the required steps. In a scenario where the test-time episode concludes either after $400$ follow steps or immediately upon a collision between the agent and the humanoid, SDA and the Baseline show a comparable collision rate ($CR$), $0.39$, and $0.38$, respectively.
@@
Heuristic Expert [36] - - 1.00 0.97 0.51 0.52 -
Baseline [36] GT 0.97±0.00 0.65±0.00 0.44±0.01 0.51±0.03 0.55±0.01*
Baseline+Proximity [9]***Code refactored and adapted from [9] to Habitat 3.0. GT 0.97±0.01 0.64±0.00 0.57±0.01 0.58±0.03 0.63±0.02
@@
Baseline [36] 0.76±0.02 0.34±0.01 0.29±0.01 0.48±0.03 0.40±0.02*
Baseline+Proximity [9] 0.85±0.02 0.41±0.02 0.37±0.01 0.58±0.02 0.41±0.01

### v2

@@ 4 Results
Section 4.1 outlines our findings on Social Navigation, along with an ablation analysis of adaptable information. Additionally, Section 4.2 offers a qualitative examination of our results, and an analysis of the role played by the latent vector $\hat{z}_{t}$. A more detailed analysis and further qualitative results can be found in the Appendix.
@@
We tested SDA on Habitat 3.0 (Puig et al., 2024), a simulation platform designed for human-robot interaction within domestic settings. This platform offers precise humanoid simulation capabilities with a focus on collaborative tasks such as Social Navigation and Social Rearrangement. It offers a vast library of avatars featuring multiple genders, body shapes, and appearances. Furthermore, it employs an oracle policy to generate movement and behavior, enabling programmable control of avatars for navigation, object interaction, and a range of other movements.
@@
The baselines we employ in our study are drawn from Habitat 3.0 (Puig et al., 2024) and consist of: (i) Heuristic Expert: a heuristic baseline equipped with access to the environment map, employing a shortest path planner to devise a route to the current location of the humanoid. The heuristic expert operates on the following principles: When the agent is beyond a distance of 1.5 meters from the humanoid, it employs a “find” behavior, utilizing a path planner to approach the humanoid. Conversely, if the humanoid is within 1.5 meters, the expert executes a backup motion to prevent collision with the humanoid. (ii) Baseline: the current SotA method (Puig et al., 2024), a recurrent neural network policy trained with DD-PPO (Wijmans et al., 2019b), operates on a “sensors-to-action” paradigm. Inputs to this policy consist of an egocentric arm and stereo depth sensors, a humanoid detector, and humanoid GPS coordinates, while the outputs are velocity commands in the robot’s local frame. Table 1 compares our model in a realistic configuration, where the humanoid GPS data are unavailable. (iii) Proximity tasks: we also adapted the Proximity Tasks defined in (Cancelli et al., 2023) and applied them to the baseline (Puig et al., 2024). These tasks were proposed for a different setup of SocialNav, where the agent acts in an environment with multiple humanoids and must navigate from point A to point B while avoiding collisions. We adapted the risk and compass proximity tasks to the SocialNav setting addressed in this article. In this context, the risk has a low value (close to 0) when the agent is within 1 to 2 meters from the humanoid and a value close to 1 when the distance is less than 1 meter or greater than 2 meters. Similarly, given the presence of only one humanoid in the environment, the compass was redefined to predict the angle between the humanoid and the agent. This adjustment aims to assist during the following phase, enabling the agent to follow the humanoid while maintaining a safe distance and staying aligned with the human. The proximity tasks are jointly trained with the policy and detached during evaluation.
@@
We used the metrics for the SocialNav task as defined in (Puig et al., 2024). Finding Success ($S$) is the ratio of the episodes where the agent located and reached the human. Finding Success Weighted by Path Steps ($SPS$) measures the optimality of the path taken by the agent wrt the optimal number of steps needed to reach the human. Following Rate ($F$) is the ratio of steps during which the robot maintains a distance of 1-2 meters from the humanoid while facing towards it relatively to the maximum possible following steps. Collision Rate ($CR$) is the ratio of the episodes that ended with the robot colliding with the humanoid. Episode Success ($ES$) measures the ratio of the episodes where the agent found the human and followed it for the required number of steps, maintaining a safe distance in the 1-2 meters range.
@@
In our work, the privileged information under consideration includes humanoid GPS (hGPS) and human trajectories (traj.). Humanoid GPS is represented in polar coordinates, a method of specifying a point’s position in a plane using two parameters: the distance from the point to the origin (radius) and the angle formed between a reference direction (typically the positive x-axis) and a line connecting the origin to the point (polar angle or azimuth). In our context, the origin is defined as the robot’s position; thus, its position is implicitly known along with that of the human. Conversely, trajectories solely consist of information derived from the human’s position within the environment.
@@ 4.1 Quantitative Results
In Table 1, we present the results obtained in Habitat 3.0 (Puig et al., 2024) for the Social Navigation task. The table is divided into three sections. The first one displays outcomes achieved by utilizing a heuristic expert endowed with extensive information, including its position, map data, and the humanoid’s position, granting it a competitive advantage over other methods. The subsequent section features models trained and tested using ground truth (GT) data (Baselines and Stage 1), thus establishing an upper limit for techniques utilizing more practical inputs feasible in real-world scenarios. Lastly, the final rows delineate results from methods conducting inference without privileged information. The models that use privileged information (GT) show that S1 has lower performance than the two Baselines, especially in $S$ and $SPS$. This is to be expected, considering that while trajectories solely depict the movement of the human in the environment, hGPS furnishes crucial details on locating the humanoid, offering insights into the distance and angle between them. In the second stage, despite the absence of human trajectory input, SDA keeps the performance level of Stage 1 by adapting to the social dynamics. Our model outperforms the baselines in the find task, increasing $S$ and $SPS$ by $6\%$.
Our approach generally improves performance in the episode success ($ES$) metric, which occurs when the agent finds the humanoid and follows it for $400$ steps. However, we emphasize that the test-time episode ends at $1500$ steps or when the agent collides with the humanoid (Puig et al., 2024), not after the $400$ follow steps. In this context, $S$, $SPS$, and $F$ metrics demonstrate how SDA, compared to the Baseline, more frequently locates the humanoid, follows a more optimal path (on average $438$ vs. $540$ steps), and follows it for a longer duration ($390$ vs. $218$ steps). Therefore, as the agent follows the humanoid longer, it has more chances to collide with it, given that the episode does not necessarily end after the required steps. In a scenario where the test-time episode concludes either after $400$ follow steps or immediately upon a collision between the agent and the humanoid, SDA and the Baseline show a comparable collision rate ($CR$), $0.39$, and $0.38$, respectively.
@@
Heuristic Expert Puig et al. (2024) - - 1.00 0.97 0.51 0.52 -
Baseline Puig et al. (2024) GT 0.97±0.00 0.65±0.00 0.44±0.01 0.51±0.03 0.55±0.01*
Baseline+Proximity Cancelli et al. (2023) [^1] Code refactored and adapted from Cancelli et al. (2023) to Habitat 3.0. GT 0.97±0.01 0.64±0.00 0.57±0.01 0.58±0.03 0.63±0.02
@@
Baseline Puig et al. (2024) 0.76±0.02 0.34±0.01 0.29±0.01 0.48±0.03 0.40±0.02*
Baseline+Proximity Cancelli et al. (2023) 0.85±0.02 0.41±0.02 0.37±0.01 0.58±0.02 0.41±0.01